# REBEL: Reinforcement Learning via Regressing Relative Rewards

### v2

@@ 2.1 Deriving REBEL: REgression to RElative REward Based RL
2.2 Extending REBEL to Stochastic MDPs
In this section, we demonstrate that REBEL can be applied to any stochastic MDP as long as we have the ability to reset to the initial states. Let $x$ denote the initial state $s_{0}$ of the MDP and $y$ represent a sequence of states and actions $(a_{0},s_{1},a_{1},...,a_{H-1},s_{H})$ where $H$ is the sequence length and $a_{h}\sim\pi_{\theta}(\cdot|s_{h})$. The probability of trajectory $y$ given $x$ is $\pi_{\theta}(y|x)=\prod_{h}\pi_{\theta}(a_{h}|s_{h})P(s_{h+1}|s_{h},a_{h})$ where $P$ is the transition probability. As in a policy gradient, observe that the transition probability cancels out since REBEL always uses probability ratios: $\pi_{\theta}(y|x)/\pi_{\theta_{t}}(y|x)=\prod_{h}\pi_{\theta}(a_{h}|s_{h})/\pi_{\theta_{t}}(a_{h}|s_{h})$. By resetting to the initial state $s_{0}$, we can generate another trajectory $y^{\prime}$ from $\pi_{\theta}$ with length $H$, leading to the following objective:
$\theta_{t+1}=\mathop{\mathrm{argmin}}_{\theta}\sum_{(x,y,y^{\prime})\in\mathcal{D}_{t}}\left(\frac{1}{\eta}\left(\sum_{h}\ln\frac{\pi_{\theta}(a_{h}|s_{h})}{\pi_{\theta_{t}}(a_{h}|s_{h})}-\sum_{h^{\prime}}\ln\frac{\pi_{\theta}(a^{\prime}_{h^{\prime}}|s^{\prime}_{h^{\prime}})}{\pi_{\theta_{t}}(a^{\prime}_{h^{\prime}}|s^{\prime}_{h^{\prime}})}\right)-\left(r(x,y)-r(x,y^{\prime})\right)\right)^{2}$
where $s_{0}=s^{\prime}_{0}=x$. We emphasize again that we only require resetting to the initial states. Note that similar to REINFORCE, the trajectory-level rewards could be high-variance, perhaps introducing the need for critic-based variance reduction (Konda and Tsitsiklis, 1999; Mnih et al., 2016; Haarnoja et al., 2018). We leave the experimental validation of REBEL in stochastic MDPs to future work.
@@ 4 Theoretical Analysis
The above theorem shows a reduction from RL to supervised learning — as long as supervised learning works (i.e., $\epsilon$ is small), then REBEL can compete against any policy $\pi^{*}$ that is covered by the base data distribution $\mu$. In the regret bound, the $1/\sqrt{T}$ comes from Mirror Descent style update, and $C_{\mu\to\pi^{*}}\epsilon$ captures the cost of distribution shift: we train our regressors under distribution $\pi_{t}$ and $\mu$, but we want the learned regressor to predict well under $\pi^{*}$. Similar to the NPG analysis from Agarwal et al. (2021), we now have a slower convergence rate $1/\sqrt{T}$, due to the fact that we have approximation error from learning. Such an agnostic regret bound — being able to compete against any policy that is covered by training distributions — is the strongest type of agnostic learning results known in the RL literature, matching the best of what has appeared in prior policy optimization work including PSDP (Bagnell et al., 2003), CPI (Kakade and Langford, 2002), NPG (Agarwal et al., 2021), and PC-PG (Agarwal et al., 2020). While in this work we use the simplest and most intuitive definition of coverage – the density ratio-based definition in Eq. 10 – extension to more general ones such as transfer error (Agarwal et al., 2020, 2021) or concentrability coefficients that incorporate the function class (e.g., Song et al. (2023)) is straightforward. We defer the proof of the above theorem and the detailed constants that we omitted in the $O$ notation to Appendix D. We include an extension of the above analysis to the general preference setting in Appendix E.
@@ 5.2 General Chat
Task. We consider a general chat scenario where $x$ is a prompt from the user and $y$ is a response. To demonstrate the robustness of our method, we apply REBEL on two entirely different sets of models and datasets. For the first set, we adapt the setting from Zhu et al. (2023), using OpenChat-3.5 (Wang et al., 2024) as the base model, Starling-RM-7B-alpha (Zhu et al., 2023) as the reward model, and the Nectar dataset (Zhu et al., 2023). This set allows us to compare REBEL to APA (Zhu et al., 2023). We compare our trained model with Starling-LM-7B-alpha. For the second set, we use Meta-Llama-3-8B-Instruct (Meta, 2024) as the base model, FsfairX-LLaMA3-RM-v0.1 (Xiong et al., 2024) as the reward model, and the UltraFeedback dataset (Cui et al., 2023).
@@
Model MT-Bench AlpacaEval 2.0
1st Turn 2nd Turn Average LC Win Rate Win Rate
Mistral-7B-Instruct-v0.2 (Jiang et al., 2023) 7.78 7.25 7.51 17.1 14.7
OpenChat-3.5 (Wang et al., 2024) 8.03 7.35 7.69 12.2 11.7
Starling-LM-7B-alpha (Zhu et al., 2023) 8.18 6.69 7.43 14.7 14.2
Meta-Llama-3-8B-Instruct (Meta, 2024) 8.31 7.89 8.10 22.9 22.6
GPT-3.5-Turbo (OpenAI, 2023) - - 7.94 18.1 9.6
Gemini Pro (Google, 2024) - - - 24.4 18.2
Claude 2 (Anthropic, 2023) - - 8.06 28.2 17.2
GPT-4 (OpenAI, 2023) - - 8.96 35.3 22.1
REBEL-OpenChat-3.5 $8.54_{(+0.51)}$ $7.58_{(+0.23)}$ $8.06_{(+0.37)}$ $17.3_{(+5.1)}$ $12.8_{(+1.1)}$
REBEL-Llama-3 $8.63_{(+0.32)}$ $7.69_{(-0.20)}$ $8.16_{(+0.06)}$ $30.1_{(+7.2)}$ $32.6_{(+10.0)}$
Table 2: Results on General Chat. The values in parentheses denote the improvement from the respective base models. The colors group the base model with the respective REBEL finetuned model.
Model MMLU GSM8K Arc Winogrande TruthfulQA HellaSwag Average
(5-shot) (5-shot) (25-shot) (5-shot) (0-shot) (10-shot)
OpenChat-3.5 (Wang et al., 2024) 63.6 68.5 64.9 80.6 47.3 84.7 68.3
Starling-LM-7B-alpha (Zhu et al., 2023) 63.4 68.0 64.9 81.1 47.3 84.8 68.2
Meta-Llama-3-8B-Instruct (Meta, 2024) 65.8 75.3 62.0 75.5 51.7 78.7 68.2
REBEL-OpenChat-3.5 $63.7_{(+0.1)}$ $68.8_{(+0.3)}$ $64.3_{(-0.6)}$ $80.4_{(-0.2)}$ $48.2_{(+0.9)}$ $85.0_{(+0.3)}$ $68.4_{(+0.1)}$
REBEL-Llama-3 $65.8_{(+0.0)}$ $75.6_{(+0.3)}$ $61.7_{(-0.3)}$ $75.8_{(+0.3)}$ $51.7_{(+0.0)}$ $78.8_{(+0.1)}$ $68.2_{(+0.0)}$
Table 3: Results on Open LLM Leaderboard. The values in parentheses denote the improvement from the respective base models. The colors group the base model with the respective REBEL finetuned model.
@@
Quality Analysis. The results between models trained with REBEL and baseline methods are shown in Table 2 and 3. For MT-Bench and AlpacaEval 2.0, under the same setup, REBEL-OpenChat-3.5 outperforms Starling-LM-7B-alpha which is trained with APA (Zhu et al., 2023) on both metrics, demonstrating the effectiveness of REBEL under chat setting and its superior performance over APA. REBEL-Llama-3 outperforms several state-of-the-art models including GPT-3.5-Turbo, Genimi Pro, and Claude 2 with only 8B parameters. For the metrics on Open LLM Leaderboard, REBEL is able to consistently enhance the performance of GSM8K and HellaSwag under different settings and maintain the overall average as the base models. Similar values on MMLU as base models indicate we preserve the basic capability of the pre-trained model during the RL fine-tuning process. We include a breakdown of MT-Bench in Appendix K.
@@ 6 Related Work
The third family of PFT techniques combines elements from the previous two: it involves running an offline algorithm iteratively, collecting on-policy preference feedback from either a supervisor model such as GPT4 (Rosset et al., 2024; Xiong et al., 2024; Guo et al., 2024) or from a preference model fit on human data (Calandriello et al., 2024). All of these approaches can be considered instantiations of the general SPO reduction proposed by Swamy et al. (2024), which itself can be thought of as a preference-based variant of DAgger (Ross et al., 2011). Recent work by Tajwar et al. (2024) confirms the empirical strength of these techniques which leverage additional online data. Our approach fits best into this family of techniques – we also iteratively update our model by solving a sequence of supervised learning problems over on-policy datasets. However, REBEL comes with several key differentiating factors from the prior work. Online versions of DPO or IPO (Xiong et al., 2024; Tajwar et al., 2024; Guo et al., 2024; Calandriello et al., 2024; Munos et al., 2023) essentially use a reward / preference model to generate binary win-loss labels while REBEL actually uses the output of the reward model as a regression target, taking advantage of this more nuanced feedback. In contrast to Rosset et al. (2024), algorithmically, REBEL does not use any online preference feedback from GPT4 nor does it require to generate a large number of responses per prompt, both of which are extremely expensive as reported by Rosset et al. (2024). Theoretically, we are able to prove policy performance bounds under a much weaker coverage condition. Unlike Mao et al. (2024) that regularize to the initial policy $\pi_{0}$ during updates, we perform conservative updates by regularizing $\pi_{t+1}$ to $\pi_{t}$. When doing the former, it is difficult to prove convergence or monotonic improvement as the current policy can just bounce around a ball centered at $\pi_{0}$, a well-known issue in the theory of approximate policy iteration (Kakade and Langford, 2002; Munos, 2003). In contrast, by incorporating the prior policy’s probabilities into our regression problem, we are able to prove stronger guarantees for REBEL. When applying REBEL to the general preference setting, our algorithm shares some similarities with the one from Wu et al. (2024). The key difference is that by using paired samples, REBEL does not need to resort to approximations of the partition function (e.g., assume the partition function is always equal to a constant). Furthermore, we can run REBEL with datasets consisting of a mixture of on-policy and off-policy data with strong guarantees, enabling hybrid training, as previously explored in the RL (Song et al., 2023; Ball et al., 2023; Zhou et al., 2023) and inverse RL (Ren et al., 2024) literature.
@@ 7 Conclusion and Future Work
There are several open questions raised by our work. The first is whether using a loss function other than square loss (e.g. log loss or cross-entropy) could lead to better performance in practice (Farebrother et al., 2024) or tighter bounds (e.g. first-order / gap-dependent) in theory (Foster and Krishnamurthy, 2021; Wang et al., 2023, 2024). The second is whether, in the general (i.e. non-utility-based) preference setting, the coverage condition assumed in our analysis is necessary – we conjecture it is. Relatedly, it would be interesting to explore whether using preference (rather than reward) models to provide supervision for REBEL replicates the performance improvements reported by Swamy et al. (2024); Munos et al. (2023). Third, while we focus primarily on the bandit setting in the preceding sections, it would be interesting to consider the more general RL setting and explore how offline datasets can be used to improve the efficiency of policy optimization via techniques like resets (Bagnell et al., 2003; Ross and Bagnell, 2014; Swamy et al., 2023; Chang et al., 2023, 2024).
@@ Acknowledgements
WS acknowledges funding from NSF IIS-2154711 and NSF CAREER 2339395; JDL acknowledges support of the NSF CCF 2002272, NSF IIS 2107304, and NSF CAREER Award 2144994. Jonathan Chang is supported by LinkedIn under the LinkedIn-Cornell Grant. Kiante Brantley is supported by NSF under grant No. 2127309 to the Computing Research Association for the CIFellows Project.
@@ Appendix D Proof of Theorem 1
where the last inequality uses Lemma 1 again. Combine things together, we can conclude that:
@@ F.1.3 Baseline Implementation Details
For supervised fine-tuning (SFT), reward modeling training, PPO, and DPO, we follow the implementation at https://github.com/vwxyzjn/summarize_from_feedback_details. For iterative dpo, we implement as follows:
@@ F.2.2 Model Details
For both OpenChat-3.5888HuggingFace Model Card: openchat/openchat_3.5 and Meta-Llama-3-8B-Instruct [^9] HuggingFace Model Card: meta-llama/Meta-Llama-3-8B-Instruct, we only train the last four layers in the model while keep other layers frozen. For both Starling-RM-7B-alpha [^10] HuggingFace Model Card: berkeley-nest/Starling-RM-7B-alpha and FsfairX-LLaMA3-RM-v0.1111111HuggingFace Model Card: sfairXC/FsfairX-LLaMA3-RM-v0.1, we directly use the reward scores without any normalizations. We filter out prompts that are longer than $1,024$ tokens ($2.3\%$) to fit the input length. Both experiments are trained on 8 H100 GPUs for four days.

### v3

@@ 4 Theoretical Analysis
The above theorem shows a reduction from RL to supervised learning — as long as supervised learning works (i.e., $\epsilon$ is small), then REBEL can compete against any policy $\pi^{*}$ that is covered by the base data distribution $\mu$. In the regret bound, the $1/\sqrt{T}$ comes from Mirror Descent style update, and $C_{\mu\to\pi^{*}}\epsilon$ captures the cost of distribution shift: we train our regressors under distribution $\pi_{t}$ and $\mu$, but we want the learned regressor to predict well under $\pi^{*}$. Similar to the NPG analysis from Agarwal et al. (2021), we now have a slower convergence rate $1/\sqrt{T}$, due to the fact that we have approximation error from learning. Such an agnostic regret bound — being able to compete against any policy that is covered by training distributions — is the strongest type of agnostic learning results known in the RL literature, matching the best of what has appeared in prior policy optimization work including PSDP (Bagnell et al., 2003), CPI (Kakade and Langford, 2002), NPG (Agarwal et al., 2021), and PC-PG (Agarwal et al., 2020). While in this work we use the simplest and most intuitive definition of coverage – the density ratio-based definition in Eq. 10 – extension to more general ones such as transfer error (Agarwal et al., 2020, 2021) or concentrability coefficients that incorporate the function class (e.g., Song et al. (2023)) is straightforward. We defer the proof of the above theorem and the detailed constants that we omitted in the $O$ notation to Appendix D.
Remark 1 (Extension to non-transitive preference).
Extending REBEL and its analysis to general non-transitive preference is straightforward, and we present the extention and its analysis in Appendix C and Appendix E, respectively.
Remark 2 (Discussion on the size of the response space $|\mathcal{Y}|$ and other design choices of the sampling distributions).
In REBEL, when sampling a pair $(y,y^{\prime})$, we in default sample $y\sim\pi_{t}$, i.e., we make sure at least one of them is an on-policy sample. This is to make sure that the training distribution at iteration $t$ covers $\pi_{t}$, which plays an essential role in avoiding a polynomial dependency on the size of the action space $|\mathcal{Y}|$. [^1] The sample complexity of the Q-NPG algorithm presented from Agarwal et al. (2019) has a polynomial dependence on the size of the action space since it samples actions uniform randomly in order to cover both $\pi_{t}$ and $\pi^{*}$. REBEL leverages that we can reset from the same context $x$, and thus directly draw two samples per context – one from $\pi_{t}$ and one from $\mu$, to cover $\pi_{t}$ and $\pi^{*}$ simultaneously. On the other hand, as long as we have some off-policy distribution $\nu_{t}$ that covers $\pi_{t}$ for all $t$, we can use it to sample $y$ and pay an additional concentrability coefficient $\max_{x,y,t}\pi_{t}(y|x)/\nu_{t}(y|x)$ in the final bound. In experiments, we test the combination of the best-of-N of $\pi_{t}$ as the base distribution $\mu$ and the worst-of-N of $\pi_{t}$ as the $\nu_{t}$. Setting $\mu$ to be the best-of-N of $\pi_{t}$ makes $\mu$ cover higher quality comparator policies. Selecting $\nu_{t}$ as the worst-of-N of $\pi_{t}$ still ensures coverage to $\pi_{t}$ while at the same time increasing the reward gap $r(x,y)-r(x,y^{\prime})$, which we find is helpful experimentally.
@@ 5.2 General Chat
Task. We consider a general chat scenario where $x$ is a prompt from the user and $y$ is a response. We adapt the setting from Zhu et al. (2023), using OpenChat-3.5 (Wang et al., 2024) as the base model, Starling-RM-7B-alpha (Zhu et al., 2023) as the reward model, and the Nectar dataset (Zhu et al., 2023). This setup enables a direct comparison between REBEL and APA (Zhu et al., 2023) which is used to train Starling-LM-7B-alpha.
@@
Method MT-Bench AlpacaEval 2.0 MMLU GSM8K Arc Winogrande TruthfulQA HellaSwag
LC Win Rate Win Rate (5-shot) (5-shot) (25-shot) (5-shot) (0-shot) (10-shot)
Base 7.69 12.2 11.7 63.6 68.5 64.9 80.6 47.3 84.7
APA 7.43 14.7 14.2 63.4 68.0 64.9 81.1 47.3 84.8
REBEL 8.06 17.3 12.8 63.7 68.8 64.3 80.4 48.2 85.0
Table 2: Results on General Chat. The best-performing method for each metric is highlighted in bold. Note that the APA result is directly obtained by evaluating the Starling-LM-7B-alpha model.
@@
Quality Analysis. The results between models trained with REBEL and baseline methods are shown in Table 2. For MT-Bench and AlpacaEval 2.0, under the same setup, REBEL outperforms APA (Zhu et al., 2023) on both metrics, demonstrating the effectiveness of REBEL under chat setting and its superior performance over APA. For the metrics on Open LLM Leaderboard, REBEL is able to enhance the performance of GSM8K and HellaSwag and maintain the overall average as the base models. Similar values on MMLU as base models indicate that we preserve the basic capability of the pre-trained model during the RL fine-tuning process. We include a breakdown of MT-Bench in Appendix K.
5.2.1 Ablation: batch size and data sampling distributions
Task. In the previous section, we sample $y$ and $y^{\prime}$ from $\pi_{t}(\cdot|x)$ and we use small batch size with $|\mathcal{D}_{t}|=32$. In this section, we investigate the alternative sampling distribution described in Remark 1. Specifically, at each iteration, we generate $5$ responses from $\pi_{t}$ for each prompt in the entire dataset (i.e., $|\mathcal{D}_{t}|$ is the size of the entire dataset), rank them based on the reward model, and set $y$ to be the best of the five responses, and $y^{\prime}$ to be the worst of the five responses. We perform $3$ iterations for this setup with Meta-Llama-3-8B-Instruct (Meta, 2024) as the base model, ArmoRM-Llama3-8B-v0.1 (Wang et al., ) as the reward model, and the UltraFeedback dataset (Cui et al., 2023). We compare REBEL with DPO which is also trained for one epoch on the entire dataset with best-of-5 as $y_{w}$ and worst-of-5 as $y_{l}$ sampled from $\pi_{0}$. In other words, the training data used for the first iteration of REBEL is the same as the one we use for DPO. [^2] Directly training DPO on the original Ultrafeedback preference dataset does not provide strong performance under AlpacaEval (e.g., the LC-winrate is around 28%, see Song et al. (2024)). So for a fair comparison to REBEL, we train DPO on the online data generated by $\pi_{0}$ and labeled by the reward model.
Evaluation. We follow the same evaluation methods as the previous section and include Arena Hard (AH) (Li et al., 2024) in our analysis.
Method MT-Bench AlpacaEval 2.0 MMLU GSM8K Arc Winogrande TruthfulQA HellaSwag AH
LC Win Rate Win Rate (5-shot) (5-shot) (25-shot) (5-shot) (0-shot) (10-shot)
Base 8.10 22.9 22.6 65.8 75.3 62.0 75.5 51.7 78.7 22.3
DPO 8.11 44.9 41.6 66.1 74.6 61.3 75.5 51.8 78.9 34.0
REBEL (iter 1) 8.13 48.3 41.8 66.3 75.8 61.7 75.9 51.8 78.7 34.5
REBEL (iter 2) 8.07 50.0 48.5 65.9 75.4 61.3 75.5 50.3 78.6 30.4
REBEL (iter 3) 8.01 49.7 48.1 66.0 75.7 61.1 75.7 49.8 78.8 30.0
Table 3: Ablation Results. In this table, REBEL uses a larger batch size (the entire dataset) with the best-of-N and worst-of-N ($N=5$) of $\pi_{t}$ as the sampling distributions for generating pairs $y,y^{\prime}$. The best-performing method for each metric is highlighted in bold. Note that DPO is trained on the online data generated by the base model and labeled by the RM.
Quality Analysis. Results in Table 3 show that REBEL can significantly improve the base model’s performance, especially on AlpacaEval 2.0 and Arena Hard. Compared to DPO, the model trained by REBEL with 1 iteration is better in almost all datasets, demonstrating the benefit of using the fine-grained reward gap in policy optimization over just the zero-one labels. In this large batch setting, we find that more iterations in general do not help performance. We conjecture that this is the issue of overfitting to the training dataset. A more diverse and larger dataset can potentially address this issue.
@@ 6 Related Work
The third family of PFT techniques combines elements from the previous two: it involves running an offline algorithm iteratively, collecting on-policy preference feedback from either a supervisor model such as GPT4 (Rosset et al., 2024; Xiong et al., 2024; Guo et al., 2024) or from a preference model fit on human data (Calandriello et al., 2024). All of these approaches can be considered instantiations of the general SPO reduction proposed by Swamy et al. (2024), which itself can be thought of as a preference-based variant of DAgger (Ross et al., 2011). Recent work by Tajwar et al. (2024) confirms the empirical strength of these techniques which leverage additional online data. Our approach fits best into this family of techniques – we also iteratively update our model by solving a sequence of supervised learning problems over on-policy datasets. However, REBEL comes with several key differentiating factors from the prior work. Online versions of DPO or IPO (Xiong et al., 2024; Tajwar et al., 2024; Guo et al., 2024; Calandriello et al., 2024; Munos et al., 2023) essentially use a reward / preference model to generate binary win-loss labels while REBEL actually uses the output of the reward model as a regression target, taking advantage of this more nuanced feedback. In contrast to Rosset et al. (2024), algorithmically, REBEL does not use any online preference feedback from GPT4 nor does it require to generate a large number of responses per prompt, both of which are extremely expensive as reported by Rosset et al. (2024). Theoretically, we are able to prove policy performance bounds under a much weaker coverage condition. Unlike Mao et al. (2024) that regularize to the initial policy $\pi_{0}$ during updates, we perform conservative updates by regularizing $\pi_{t+1}$ to $\pi_{t}$. When doing the former, it is difficult to prove convergence or monotonic improvement as the current policy can just bounce around a ball centered at $\pi_{0}$, a well-known issue in the theory of approximate policy iteration (Kakade and Langford, 2002; Munos, 2003). In contrast, by incorporating the prior policy’s probabilities into our regression problem, we are able to prove stronger guarantees for REBEL. When applying REBEL to the general preference setting, our algorithm shares some similarities with the concurrent work of Wu et al. (2024), which also leverages the DPO reparameterization trick to cleanly implement an Online Mirror Descent approach to computing a minimax winner via self-play. The key difference is that REBEL uses paired responses, while the algorithm of Wu et al. (2024) does not. Using paired responses, REBEL is able to cancel out the partition function and therefore does not need to resort to heuristic approximations of it (specifically, assuming the partition function is always equal to a constant). Furthermore, we can run REBEL with datasets consisting of a mixture of on-policy and off-policy data with strong guarantees, enabling hybrid training, as previously explored in the RL (Song et al., 2023; Ball et al., 2023; Zhou et al., 2023) and inverse RL (Ren et al., 2024) literature.
@@ 7 Conclusion and Future Work
There are several open questions raised by our work. The first is whether using a loss function other than square loss (e.g. log loss or cross-entropy) could lead to better performance in practice (Farebrother et al., 2024) or tighter bounds (e.g. first-order / gap-dependent) in theory (Foster and Krishnamurthy, 2021; Wang et al., 2023, 2024). The second is whether, in the general (i.e. non-utility-based) preference setting, the coverage condition assumed in our analysis is necessary – we conjecture it is. Relatedly, it would be interesting to explore whether using preference (rather than reward) models to provide supervision for REBEL replicates the performance improvements reported by Swamy et al. (2024); Munos et al. (2023); Calandriello et al. (2024). Third, while we focus primarily on the bandit setting in the preceding sections, it would be interesting to consider the more general RL setting and explore how offline datasets can be used to improve the efficiency of policy optimization via techniques like resets (Bagnell et al., 2003; Ross and Bagnell, 2014; Swamy et al., 2023; Chang et al., 2023, 2024; Ren et al., 2024; Dice et al., 2024).
@@ Acknowledgements
WS acknowledges funding from NSF IIS-2154711, NSF CAREER 2339395, DARPA LANCER: LeArning Network CybERagents, and Cornell Infosys Collaboration; JDL acknowledges support of the NSF CCF 2002272, NSF IIS 2107304, and NSF CAREER Award 2144994. Jonathan Chang is supported by LinkedIn under the LinkedIn-Cornell Grant. Kiante Brantley is supported by NSF under grant No. 2127309 to the Computing Research Association for the CIFellows Project. GKS is supported by his family and friends.
@@ Appendix D Proof of Theorem 1
where the last inequality uses Lemma 1 again. This step relies on the fact that one of the samples is always on-policy, i.e., from $\pi_{t}$.
Combine things together, we can conclude that:
@@ F.1.3 Baseline Implementation Details
For supervised fine-tuning (SFT), reward modeling training, PPO, and DPO, we follow the implementation at https://github.com/vwxyzjn/summarize_from_feedback_details (Huang et al., 2024). For iterative dpo, we implement as follows:
@@ F.2.2 Model Details
For OpenChat-3.5101010HuggingFace Model Card: openchat/openchat_3.5, we only train the last four layers and keep other layers frozen. For Meta-Llama-3-8B-Instruct [^11] HuggingFace Model Card: meta-llama/Meta-Llama-3-8B-Instruct, we perform full-parameter training. For Starling-RM-7B-alpha [^12] HuggingFace Model Card: berkeley-nest/Starling-RM-7B-alpha and ArmoRM-Llama3-8B-v0.1131313HuggingFace Model Card: RLHFlow/ArmoRM-Llama3-8B-v0.1, we directly use the reward scores without any normalizations. We filter out prompts that are longer than $1,024$ tokens ($2.3\%$) to fit the input length. OpenChat-3.5 is trained for four days, and Meta-Llama-3-8B-Instruct is train for one day on 8 H100 GPUs.